# AlignDiT: Multimodal Aligned Diffusion Transformer for Synchronized Speech Generation

### v1

@@ Abstract
Figure 1: AlignDiT aims to generate natural speech that is (1) accurately aligned with the given text, (2) temporally synchronized with the input video, and (3) acoustically consistent with the reference speech. This versatile system has a broad range of applications, including (a) Automated dialogue replacement, (b) Video-to-Speech synthesis, and (c) Visual Forced Alignment.
1 Introduction
@@
Human communication involves the exchange of information through spoken language, which serves as a fundamental means of interaction between individuals [49, 27, 18]. It naturally generates multimodal signals that convey both linguistic content and paralinguistic cues. Among them, audio, video, and text are three key modalities widely utilized in speech and language processing. The audio modality delivers the sound signal that carries phonetic content, prosodic features such as intonation and rhythm, as well as speaker-specific characteristics. The video modality captures lip movements, facial expressions, and non-verbal signals that are synchronized with the audio. Text, while not directly observable in the physical world, serves as a symbolic and discrete representation of the linguistic content, monotonically aligned with audio and video over time [57, 62]. Together, these modalities complement one another, offering a comprehensive view of human verbal communication.
@@
Cross-modal generation tasks, where one modality is synthesized or inferred from others, have been extensively studied to support accessibility and multimodal interaction. Typical examples include text-to-speech synthesis (TTS) [63, 7, 8, 50], automatic speech recognition (ASR) [64, 4, 22, 61], lip reading [66, 45, 46, 75], video-to-speech synthesis [51, 74, 32, 12], and talking face generation [59, 55, 71]. While many of these tasks focus on converting a single input modality into another, incorporating multiple input modalities can lead to more accurate and robust generation, as different modalities provide complementary information. For example, audio-visual speech recognition (AVSR) [1, 28, 11] improves transcription accuracy by leveraging both audio and video inputs, especially under noisy conditions.
@@
This motivates generating speech audio conditioned on both video and text inputs, where visual cues from lip movements and explicit linguistic content complement each other. A representative application of this task is Automated Dialogue Replacement (ADR) [14, 15], widely used in film and television post-production. During this process, actors’ lines are re-recorded to improve audio quality while ensuring synchronization with the video, especially when original recordings are affected by background noise or challenging recording conditions. Building on recent advances in deep learning based speech generative models, prior studies have explored generating speech from silent video, transcriptions, and short reference audio clips. Despite their potential, existing multimodal-to-speech generation approaches face three key challenges. First, they often struggle to generate natural and intelligible speech due to limited modeling capacity and reliance on limited datasets. Second, they are not robust when one or more input modalities are missing or corrupted, as they lack mechanisms to adjust the importance of each modality. Third, many methods depend on external forced aligners or duration predictors for synchronization, increasing supervision costs and risks propagating alignment errors.
@@
Figure 2: Various methods for conditioning multimodal inputs to DiT blocks: (a) channel-wise concatenation of text, reference speech, and visual features; (b) providing the text as a prefix; and (c) conditioning text inputs through cross-attention. In (a), $\phi$ represents filler tokens.
2 Related Works
2.1 Multimodal Speech Tasks
@@
Automated dialogue replacement (ADR). Early efforts in ADR, also known as automated video dubbing, framed it as a multimodal text-to-speech (TTS) problem. Neural Dubber [30] pioneered this direction by generating speech from text conditioned on video lip movements. Subsequent work, VisualTTS [44], integrated visual information more explicitly into the TTS pipeline by introducing a textual-visual attention mechanism to learn alignments between phonemes and lip frames, as well as visual feature fusion during acoustic decoding. VDTTS [23] extended this idea to unconstrained, multi-speaker settings, predicting finer prosodic elements for a more natural dubbing. More recently, HPMDubbing [14] presented a unified architecture employing hierarchical prosody modeling. It extracts visual features at the lip, face, and scene levels to control different aspects of speech (timing, energy/pitch, and global emotion), synthesizing natural and emotional speech and setting a strong state-of-the-art benchmark for ADR. StyleDubber [15] further focused on speaker style and pronunciation habits by introducing two-level style learning (phoneme-level and utterance-level) using reference audio. Both HPMDubbing and StyleDubber require explicit alignment through a forced aligner during training to ensure accurate synchronization between video and speech.
@@
Video-to-speech. Distinct from ADR, video-to-speech generates speech directly from silent videos without transcripts. Initial methods approached this task using CNN-based [19, 37] or sequence-to-sequence models [58] to capture lip movements and generate speech. Recent approaches have further improved speech generation quality by incorporating advanced techniques, such as GANs [34, 52], normalizing flows [24, 32], and diffusion models [9, 74]. Additionally, to capture speaker characteristics, most works [10, 51] utilize speaker embeddings derived from reference audio. As obtaining reference audio during inference is not always feasible, several studies [9, 74, 32, 33] extract speaker information directly from the given video to facilitate speaker-aware speech synthesis.
@@
Visual forced alignment (VFA). VFA is a task that identifies the timeline—specifically, the start and end times—for each word or phoneme in silent videos based on corresponding textual content. VFA requires accurately aligning lip movements with text. One approach employs visual keyword spotting [60, 53], a technique identifying durations of individual words. Repeating this across an entire video achieves visual forced alignment, but involves substantial computational demands and potential inaccuracies due to overlaps between adjacent words. Another common approach [38] utilizes Connectionist Temporal Classification (CTC) [21], frequently employed in visual speech recognition (VSR) for video-text alignment. DVFA [35] represents the first method explicitly designed for VFA, leveraging a multimodal attention mechanism to effectively align textual transcriptions with corresponding lip movements.
@@
2.2 Text-to-speech (TTS) Synthesis
@@
Recent research in Text-to-Speech (TTS) synthesis has achieved significant advancements. One notable approach, autoregressive (AR)-based TTS models [69, 7, 31], combines powerful speech tokenizers [16, 77] with next-token prediction language modeling and has demonstrated promising results. Despite their high quality, non-autoregressive (NAR) models benefit from fast inference due to parallel processing, effectively balancing quality and latency. Specifically, diffusion models [26, 67] have significantly contributed to the success of current NAR approaches. For example, Matcha-TTS [50] adopts Flow Matching with Optimal Transport paths (FM-OT) [41] for training and relies on a phoneme-level duration model for speech synthesis. DiTTo-TTS [40] improves alignment by utilizing a Diffusion Transformer (DiT) [56] with cross-attention conditioned on encoded text from a pretrained language model. E2 TTS [20] removes phoneme and duration predictors, directly using characters padded with filler tokens to match the length of mel spectrograms. Additionally, F5-TTS [8] enhances text-speech alignment by integrating ConvNeXt V2 [70] into a diffusion transformer framework. In this work, we aim to extend NAR TTS models, specifically DiT, to multimodal TTS synthesis. By leveraging classifier-free guidance and carefully designed conditioning, our approach flexibly handles varying inputs—including text-only, video-only, and multimodal scenarios—for high-quality speech synthesis.
3 AlignDiT
@@
Our goal is to generate a speech waveform that matches the lip movements in a video, conveys the content of a given text script, and resembles the voice characteristics of a target speaker indicated by reference speech. We employ an in-context learning-based speech synthesis approach to handle multimodal inputs. For text-to-speech synthesis task, various methods [20, 8, 73, 40] have explored diverse strategies for conditioning text inputs. However, to the best of our knowledge, no existing method simultaneously handles video input alongside text and reference speech to achieve multimodal alignment for speech synthesis within in-context learning-based generative models. We aim to explore various settings and identify the most suitable method for multimodal alignment, capable of flexibly handling both unimodal and multimodal inputs to synthesize high-quality and accurate speech.
3.1 Model Architecture
@@
We utilize the Diffusion Transformer (DiT) [56] for multimodal-to-speech generation, as it has been shown to be effective for speech generation task. Following prior works about text-to-speech [39, 8], we train the model using a flow matching objective and generate mel-spectrogram through an iterative inference process starting from random noise. The model consists of multiple blocks, each following the standard Transformer [68] architecture with self-attention, along with additional parameters for conditioning diffusion timestep information. For multimodal-to-speech generation, we guide the generative process using a fused multimodal representation as a condition, allowing the model to progressively refine the noisy mel-spectrogram into a representation that aligns with the given multimodal inputs.
3.2 Multimodal Conditioning
@@
Audio-Video Fusion. Audio and video are naturally synchronized modalities. When a person speaks, their lip movements correspond closely to the acoustic features at that same moment, making frame-by-frame fusion straightforward. However, since the frame rates of the two modalities typically differ, we extract and encode each modality to have a common temporal resolution. Specifically, audio is converted into mel‐spectrogram sequence $h_{audio}$ at 100 fps. For video, we first extract a sequence of video features $h_{video}$ using a pretrained video encoder specialized in lip motion. To match the frame rate, we upsample the 25 fps video features to 100 fps via lightweight transposed convolutional layers. We then apply Conformer [22] encoder for better capturing contextual information. Once aligned temporally, the audio and video features are concatenated channel‐wise to form a unified multimodal representation, which effectively encodes when and how to speak.
@@
$h_{av}=[(1-M)\odot h_{audio};M\odot h_{video}]\in\mathbb{R}^{T\times 2D},$ (1)
@@
Audio-Video-Text Fusion. While the audio‐video streams are inherently time-aligned, text is only monotonically aligned with them and lacks strict frame-level synchronization. Unlike previous methods that rely on text token-level duration information from external forced aligners [48] and duration predictors [63], we aim to train the generative model fuse the modalities naturally, without explicit duration constraints. This simplifies the data preprocessing pipeline, making it easier to scale the dataset, and avoids potential biases introduced by forced alignments, thereby facilitating a more natural audio-video-text alignment. To investigate feasible modality fusion methods in the DiT blocks, we explore three conditioning strategies as illustrated in Fig. 2. In all cases, for text encoding, the character sequence is first embedded through a lookup table and then refined by a convolutional encoder to be $h_{text}$ which has the length of $L$.
(a) Early Fusion & Self‐Attention: A naive approach for fusion is concatenating all conditioning modalities along the channel axis. To match the total length, we add filler tokens at the end of the text sequence inspired by E2 TTS [20]. After this early fusion, the self-attention layers in DiT blocks learn alignment naturally to predict the masked part of mel-spectrogram. This can be expressed as follows:
$h^{\prime}=[h_{av};h_{text}]\in\mathbb{R}^{T\times 3D},$ (2)
$h=h^{\prime}W_{a}+b_{a}\in\mathbb{R}^{T\times D},$ (3)
@@
(b) Prefix & Self‐Attention: Another approach for fusion is frame-wise concatenation between text feature and audio-video feature, treating the text as a prefix. This leverages in-context conditioning [69, 73], allowing the text prefix to guide the model within the self-attention layers of DiT blocks. After the model outputs the full sequence, the prefix part corresponding to the text length is discarded, yielding the predicted mel-spectrogram. This can be formulated as follows:
$h^{\prime}=\text{Concat}(h_{text},h_{av})\in\mathbb{R}^{(L+T)\times 2D},$ (4)
$h=h^{\prime}W_{b}+b_{b}\in\mathbb{R}^{(L+T)\times D},$ (5)
@@
$h=\text{MHCA}(h_{av}W_{Q},h_{text}W_{K},h_{text}W_{V})\in\mathbb{R}^{T\times D},$ (6)
@@
3.3 Training Objective
Based on the fused multimodal representations, we train our AlignDiT with multi-task learning. First, we adopt the Conditional Flow Matching (CFM) training objective, which has been proven its effectiveness in generating high-quality data samples in an efficient manner [41]. The CFM seeks to match a probability path $p_{t}$ from a tractable distribution $p_{0}$ to $p_{1}$ approximating the target distribution. Given the fused multimodal representation $h$ and a noisy mel-spectrogram $x_{t}$ ($t\in[0,1]$), AlignDiT is trained to regress the vector field $u_{t}$ with CFM objectives:
$\mathcal{L}_{CFM}=\mathbb{E}_{t,p_{t}}\|v_{t}(x_{t}|h,\theta)-u_{t}(x_{t})\|^{2},$ (7)

### v2

@@ Abstract.
Diffusion Transformer, Speech Generation, Multimodal Alignment, Automated Dialogue Replacement
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754577isbn: 979-8-4007-2035-2/2025/10submissionid: mfp7205ccs: Information systems Multimedia content creation
@@
Figure 1. AlignDiT aims to generate natural speech that is (1) accurately aligned with the given text, (2) temporally synchronized with the input video, and (3) acoustically consistent with the reference speech. This versatile system has a broad range of applications, including (a) Automated dialogue replacement, (b) Video-to-Speech synthesis, and (c) Visual Forced Alignment.
1. Introduction
@@
Human communication involves the exchange of information through spoken language, which serves as a fundamental means of interaction between individuals (McGurk and MacDonald, 1976; Holler and Levinson, 2019; Drijvers and Holler, 2023). It naturally generates multimodal signals that convey both linguistic content and paralinguistic cues. Among them, audio, video, and text are three key modalities widely utilized in speech and language processing. The audio modality delivers the sound signal that carries phonetic content, prosodic features such as intonation and rhythm, as well as speaker-specific characteristics. The video modality captures lip movements, facial expressions, and non-verbal signals that are synchronized with the audio. Text, while not directly observable in the physical world, serves as a symbolic and discrete representation of the linguistic content, monotonically aligned with audio and video over time (Petridis et al., 2018; Ren et al., 2019). Together, these modalities complement one another, offering a comprehensive view of human verbal communication.
@@
Cross-modal generation tasks, where one modality is synthesized or inferred from others, have been extensively studied to support accessibility and multimodal interaction. Typical examples include text-to-speech synthesis (TTS) (Ren et al., 2021; Chen et al., 2024, 2025; Mehta et al., 2024), automatic speech recognition (ASR) (Schneider et al., 2019; Baevski et al., 2020; Gulati et al., 2020; Radford et al., 2023), lip reading (Son Chung et al., 2017; Ma et al., 2021, 2023; Yeo et al., 2024), video-to-speech synthesis (Mira et al., 2022a; Yemini et al., 2024; Kim et al., 2024; Choi et al., 2025a), and talking face generation (Prajwal et al., 2020b; Park et al., 2022; Xu et al., 2024). While many of these tasks focus on converting a single input modality into another, incorporating multiple input modalities can lead to more accurate and robust generation, as different modalities provide complementary information. For example, audio-visual speech recognition (AVSR) (Afouras et al., 2018b; Hong et al., 2023; Choi et al., 2024) improves transcription accuracy by leveraging both audio and video inputs, especially under noisy conditions.
@@
This motivates generating speech audio conditioned on both video and text inputs, where visual cues from lip movements and explicit linguistic content complement each other. A representative application of this task is Automated Dialogue Replacement (ADR) (Cong et al., 2023, 2024), widely used in film and television post-production. During this process, actors’ lines are re-recorded to improve audio quality while ensuring synchronization with the video, especially when original recordings are affected by background noise or challenging recording conditions. Building on recent advances in deep learning based speech generative models, prior studies have explored generating speech from silent video, transcriptions, and short reference audio clips. Despite their potential, existing multimodal-to-speech generation approaches face three key challenges. First, they often struggle to generate natural and intelligible speech due to limited modeling capacity and reliance on limited datasets. Second, they are not robust when one or more input modalities are missing or corrupted, as they lack mechanisms to adjust the importance of each modality. Third, many methods depend on external forced aligners or duration predictors for synchronization, increasing supervision costs and risks propagating alignment errors.
@@
Figure 2. Various methods for conditioning multimodal inputs to DiT blocks: (a) channel-wise concatenation of text, reference speech, and visual features; (b) providing the text as a prefix; and (c) conditioning text inputs through cross-attention. In (a), $\phi$ represents filler tokens.
2. Related Works
2.1. Multimodal Speech Tasks
@@
Automated dialogue replacement (ADR). Early efforts in ADR, also known as automated video dubbing, framed it as a multimodal text-to-speech (TTS) problem. Neural Dubber (Hu et al., 2021) pioneered this direction by generating speech from text conditioned on video lip movements. Subsequent work, VisualTTS (Lu et al., 2022), integrated visual information more explicitly into the TTS pipeline by introducing a textual-visual attention mechanism to learn alignments between phonemes and lip frames, as well as visual feature fusion during acoustic decoding. VDTTS (Hassid et al., 2022) extended this idea to unconstrained, multi-speaker settings, predicting finer prosodic elements for a more natural dubbing. More recently, HPMDubbing (Cong et al., 2023) presented a unified architecture employing hierarchical prosody modeling. It extracts visual features at the lip, face, and scene levels to control different aspects of speech (timing, energy/pitch, and global emotion), synthesizing natural and emotional speech and setting a strong state-of-the-art benchmark for ADR. StyleDubber (Cong et al., 2024) further focused on speaker style and pronunciation habits by introducing two-level style learning (phoneme-level and utterance-level) using reference audio. Both HPMDubbing and StyleDubber require explicit alignment through a forced aligner during training to ensure accurate synchronization between video and speech.
@@
Video-to-speech. Distinct from ADR, video-to-speech generates speech directly from silent videos without transcripts. Initial methods approached this task using CNN-based (Ephrat and Peleg, 2017; Kumar et al., 2019) or sequence-to-sequence models (Prajwal et al., 2020a) to capture lip movements and generate speech. Recent approaches have further improved speech generation quality by incorporating advanced techniques, such as GANs (Kim et al., 2021; Mira et al., 2022b), normalizing flows (He et al., 2022; Kim et al., 2024), and diffusion models (Choi et al., 2023a; Yemini et al., 2024). Additionally, to capture speaker characteristics, most works (Choi et al., 2023b; Mira et al., 2022a) utilize speaker embeddings derived from reference audio. As obtaining reference audio during inference is not always feasible, several studies (Choi et al., 2023a; Yemini et al., 2024; Kim et al., 2024, 2025) extract speaker information directly from the given video to facilitate speaker-aware speech synthesis.
@@
Visual forced alignment (VFA). VFA is a task that identifies the timeline—specifically, the start and end times—for each word or phoneme in silent videos based on corresponding textual content. VFA requires accurately aligning lip movements with text. One approach employs visual keyword spotting (Prajwal et al., 2021; Momeni et al., 2020), a technique identifying durations of individual words. Repeating this across an entire video achieves visual forced alignment, but involves substantial computational demands and potential inaccuracies due to overlaps between adjacent words. Another common approach (Kürzinger et al., 2020) utilizes Connectionist Temporal Classification (CTC) (Graves et al., 2006), frequently employed in visual speech recognition (VSR) for video-text alignment. DVFA (Kim et al., 2023) represents the first method explicitly designed for VFA, leveraging a multimodal attention mechanism to effectively align textual transcriptions with corresponding lip movements.
@@
2.2. Text-to-speech (TTS) Synthesis
@@
Recent research in Text-to-Speech (TTS) synthesis has achieved significant advancements. One notable approach, autoregressive (AR)-based TTS models (Wang et al., 2023; Chen et al., 2024; Kharitonov et al., 2023), combines powerful speech tokenizers (Défossez et al., 2023; Zeghidour et al., 2021) with next-token prediction language modeling and has demonstrated promising results. Despite their high quality, non-autoregressive (NAR) models benefit from fast inference due to parallel processing, effectively balancing quality and latency. Specifically, diffusion models (Ho et al., 2020; Song et al., 2020) have significantly contributed to the success of current NAR approaches. For example, Matcha-TTS (Mehta et al., 2024) adopts conditional flow matching with optimal transport paths (OT-CFM) (Lipman et al., 2023) for training and relies on a phoneme-level duration model for speech synthesis. DiTTo-TTS (Lee et al., 2025) improves alignment by utilizing a Diffusion Transformer (DiT) (Peebles and Xie, 2023) with cross-attention conditioned on encoded text from a pretrained language model. E2 TTS (Eskimez et al., 2024) removes phoneme and duration predictors, directly using characters padded with filler tokens to match the length of mel spectrograms. Additionally, F5-TTS (Chen et al., 2025) enhances text-speech alignment by integrating ConvNeXt V2 (Woo et al., 2023) into a diffusion transformer framework. In this work, we aim to extend NAR TTS models, specifically DiT, to multimodal TTS synthesis. By leveraging classifier-free guidance and carefully designed conditioning, our approach flexibly handles varying inputs—including text-only, video-only, and multimodal scenarios—for high-quality speech synthesis.
3. AlignDiT
@@
Our goal is to generate a speech waveform that matches the lip movements in a video, conveys the content of a given text script, and resembles the voice characteristics of a target speaker indicated by reference speech. We employ an in-context learning-based speech synthesis approach to handle multimodal inputs. For text-to-speech synthesis task, various methods (Eskimez et al., 2024; Chen et al., 2025; Yang et al., 2024; Lee et al., 2025) have explored diverse strategies for conditioning text inputs. However, to the best of our knowledge, no existing method simultaneously handles video input alongside text and reference speech to achieve multimodal alignment for speech synthesis within in-context learning-based generative models. We aim to explore various settings and identify the most suitable method for multimodal alignment, capable of flexibly handling both unimodal and multimodal inputs to synthesize high-quality and accurate speech.
3.1. Model Architecture
@@
We utilize the Diffusion Transformer (DiT) (Peebles and Xie, 2023) for multimodal-to-speech generation, as it has been shown to be effective for speech generation task. Following prior works about text-to-speech (Le et al., 2023; Chen et al., 2025), we train the model using a flow matching objective and generate mel-spectrogram through an iterative inference process starting from random noise. The model consists of multiple blocks, each following the standard Transformer (Vaswani et al., 2017) architecture with self-attention, along with additional parameters for conditioning diffusion timestep information. For multimodal-to-speech generation, we guide the generative process using a fused multimodal representation as a condition, allowing the model to progressively refine the noisy mel-spectrogram into a representation that aligns with the given multimodal inputs.
3.2. Multimodal Conditioning
@@
Audio-Video Fusion. Audio and video are naturally synchronized modalities. When a person speaks, their lip movements correspond closely to the acoustic features at that same moment, making frame-by-frame fusion straightforward. However, since the frame rates of the two modalities typically differ, we extract and encode each modality to have a common temporal resolution. Specifically, audio is converted into mel‐spectrogram sequence $h_{audio}$ at 100 fps. For video, we first extract a sequence of video features $h_{video}$ using a pretrained video encoder specialized in lip motion. To match the frame rate, we upsample the 25 fps video features to 100 fps via lightweight transposed convolutional layers. We then apply Conformer (Gulati et al., 2020) encoder for better capturing contextual information. Once aligned temporally, the audio and video features are concatenated channel‐wise to form a unified multimodal representation, which effectively encodes when and how to speak.
@@
(1) $h_{av}=[(1-M)\odot h_{audio};M\odot h_{video}]\in\mathbb{R}^{T\times 2D},$
@@
Audio-Video-Text Fusion. While the audio‐video streams are inherently time-aligned, text is only monotonically aligned with them and lacks strict frame-level synchronization. Unlike previous methods that rely on text token-level duration information from external forced aligners (McAuliffe et al., 2017) and duration predictors (Ren et al., 2021), we aim to train the generative model fuse the modalities naturally, without explicit duration constraints. This simplifies the data preprocessing pipeline, making it easier to scale the dataset, and avoids potential biases introduced by forced alignments, thereby facilitating a more natural audio-video-text alignment. To investigate feasible modality fusion methods in the DiT blocks, we explore three conditioning strategies as illustrated in Fig. 2. In all cases, for text encoding, the character sequence is first embedded through a lookup table and then refined by a convolutional encoder to be $h_{text}$ which has the length of $L$.
(a) Early Fusion & Self‐Attention: A naive approach for fusion is concatenating all conditioning modalities along the channel axis. To match the total length, we add filler tokens at the end of the text sequence inspired by E2 TTS (Eskimez et al., 2024). After this early fusion, the self-attention layers in DiT blocks learn alignment naturally to predict the masked part of mel-spectrogram. This can be expressed as follows:
(2) $h^{\prime}=[h_{av};h_{text}]\in\mathbb{R}^{T\times 3D},$
(3) $h=h^{\prime}W_{a}+b_{a}\in\mathbb{R}^{T\times D},$
@@
(b) Prefix & Self‐Attention: Another approach for fusion is frame-wise concatenation between text feature and audio-video feature, treating the text as a prefix. This leverages in-context conditioning (Wang et al., 2023; Yang et al., 2024), allowing the text prefix to guide the model within the self-attention layers of DiT blocks. After the model outputs the full sequence, the prefix part corresponding to the text length is discarded, yielding the predicted mel-spectrogram. This can be formulated as follows:
(4) $h^{\prime}=\text{Concat}(h_{text},h_{av})\in\mathbb{R}^{(L+T)\times 2D},$
(5) $h=h^{\prime}W_{b}+b_{b}\in\mathbb{R}^{(L+T)\times D},$
@@
(6) $h=\text{MHCA}(h_{av}W_{Q},h_{text}W_{K},h_{text}W_{V})\in\mathbb{R}^{T\times D},$
@@
3.3. Training Objective
Based on the fused multimodal representations, we train our AlignDiT with multi-task learning. First, we adopt the conditional flow matching (CFM) training objective, which has been proven its effectiveness in generating high-quality data samples in an efficient manner (Lipman et al., 2023). The CFM seeks to match a probability path $p_{t}$ from a tractable distribution $p_{0}$ to $p_{1}$ approximating the target distribution. Given the fused multimodal representation $h$ and a noisy mel-spectrogram $x_{t}$ ($t\in[0,1]$), AlignDiT is trained to regress the vector field $u_{t}$ with CFM objectives:
(7) $\mathcal{L}_{CFM}=\mathbb{E}_{t,p_{t}}\|v_{t}(x_{t}|h,\theta)-u_{t}(x_{t})\|^{2},$